OpenStudy (anonymous):

the population of a city is 2,500 . if the number of males is 240 more then the number of females, how many males and females are there in the city

OpenStudy (narissa):

try half of 2500 then add 240

Directrix (directrix):

Let F be the number of females. Let M be the number of males. M = F + 240 M + F = 2500 F + 240 + F = 2500 @kyleighA Solve for F. Then, using M = F + 240, substitute for F and calculate the number of M.